# WaveRoRA: Wavelet Rotary Route Attention for Multivariate Time Series Forecasting

### v1

@@ IV-C Multi-scale Wave Embedding
The input MTS data is first decomposed to $J$ high-pass coefficients $y_{h}^{(j)}\in\mathbb{R}^{M\times L^{(j)}}$ and 1 low-pass coefficients $y_{l}^{(J)}\in\mathbb{R}^{M\times L^{(J)}}$ with $J$-level DWT, where $j\in\{1,2,...,J\}$. Each coefficient series is passed through a corresponding wave embedding layer defined in Eq 6, where $\mathcal{C}=\{yh^{(1)},yh^{(2)}...,yh^{(J)},yl^{(J)}\}$ and $\mathrm{WaveEmbedding}$ is a linear projector. Specifically, $\mathrm{WaveEmbedding}_{j}$ contains a weight matrix $W_{j}$ and a bias matrix $b_{j}$. Note that $W_{j}\in\mathbb{R}^{D\times L^{(j)}}$ for $j\in\{1,2,...,J\}$ and $W_{J+1}\in\mathbb{R}^{D\times L^{(J)}}$, while $b_{j}\in\mathbb{R}^{D\times 1}$. The multi-scale wave embedding layers unify the feature dimensions of wavelet coefficients obtained from the multi-layer DWT. What’s more, the embedding process facilitates in capturing the temporal patterns within the time series.
@@
IV-D Rotary Route Attention
The RoRA is proposed to capture inter-series dependencies in the wavelet domain. We first transpose $\mathbb{E}\in\mathbb{R}^{(J+1)\times M\times D}$ into a sequence $\mathbb{S}\in\mathbb{R}^{M\times(J+1)\times D}$ consisting of individual variables. Specifically, we concatenate the wavelet embeddings of each variable to form $\mathbb{S}_{m}\in\mathbb{R}^{(J+1)\times D}$, where $m\in\{1,2,...,M\}$. To simplify the representation, we use $D^{\prime}$ as the hidden dimension of each sequence token instead of $(J+1)\times D$ and denote $I\in\mathbb{R}^{M\times D^{\prime}}$ as the input of RoRA.
$R_{\Theta}^{d}=\begin{pmatrix}\cos{\theta_{1}}&-\sin{\theta_{1}}&0&0&\cdots&0&0\\ \sin{\theta_{1}}&\cos{\theta_{1}}&0&0&\cdots&0&0\\ 0&0&\cos{\theta_{2}}&-\sin{\theta_{2}}&\cdots&0&0\\ 0&0&\sin{\theta_{2}}&\cos{\theta_{2}}&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&0&\cdots&\cos{\theta_{\frac{D^{\prime}}{2}}}&-\sin{\theta_{\frac{D^{\prime}}{2}}}\\ 0&0&0&0&\cdots&\sin{\theta_{\frac{D^{\prime}}{2}}}&\cos{\theta_{\frac{D^{\prime}}{2}}}\end{pmatrix}$ (7)
@@ IV-D Rotary Route Attention
We first utilizes a set of linear projectors to generate query, key and value matrices, represented as $Q\in\mathbb{R}^{M\times D^{\prime}}$, $K\in\mathbb{R}^{M\times D^{\prime}}$ and $V\in\mathbb{R}^{M\times D^{\prime}}$. Subsequently, a rotary matrix $\mathcal{R}_{\Theta}^{D^{\prime}}$ is applied to $Q$ and $K$ to inject relative positional information[34]. Specifically, $\mathcal{R}_{\Theta}^{D^{\prime}}$ consists of $M$ sub rotary matrices which can be defined in Eq 7. We then randomly initialize a series of routing tokens $R\in\mathbb{R}^{r\times D^{\prime}}$ with a small number of $r$. The route attention can be represented in Eq 8:
$Q^{\mathcal{R}}$ $=\mathcal{R}_{\Theta}^{D^{\prime}}Q,~{}K^{\mathcal{R}}=\mathcal{R}_{\Theta}^{D^{\prime}}K$ (8)
$V_{R}$ $=\mathrm{Attention}(R,K^{\mathcal{R}},V)$
$V_{O}$ $=\mathrm{Attention}(Q^{\mathcal{R}},R,V_{R})+\mathrm{Linear}_{skip}(V)$
where $\mathrm{Attention}$ is the function of Softmax attention and $\mathrm{Linear}_{skip}$ refers to the skip connection layer which promotes faster model convergence. The routing tokens $R$ is initially viewed as the query for $K^{\mathcal{R}}$ and $V$, aggregating routing features from all keys and values. Afterward, $R$ and $V_{R}$ will serve as the key matrix and value matrix respectively, enabling the transfer of information from the routing features to each query token in $Q$. In our implementations, $R$ is expected to prune redundant relationships between low-correlation variables while preserving interaction information among high-correlation variables. In addition, we set $r\ll M$ as a fixed number to achieve linear complexity of $O(Mrd)$ relative to the variable number $M$. The complete operation of RoRA mechanism is defined as Eq 9, where $\mathrm{RoRA}$ corresponds to Eq 8 and $\odot$ refers to Hadamard product. The activation function $\sigma$ is set as $\mathrm{SiLU}$[36], generating $G\in\mathbb{R}^{M\times D^{\prime}}$ which serves as a gating function.
$Q$ $=\mathrm{Linear}_{Q}(I)$ (9)
$K$ $=\mathrm{Linear}_{K}(I)$
$V$ $=\mathrm{Linear}_{V}(I)$
$V_{O}$ $=\mathrm{RoRA}(Q,K,V)$
$G$ $=\sigma(\mathrm{Linear}(I))$
$O$ $=\mathrm{Linear}(V_{O}\odot G)$
IV-E Wave-wise Normalization
@@
Considering that the wavelet coefficients of different decomposition layers correspond to different frequency intervals, we transpose the output of RoRA and apply wave-wise normalization for wave embeddings of each decomposition layer. The process corresponds to ”WaveNorm” in Fig 2. Specifically, we calculate $\mathbb{O}=\mathrm{Transpose}(O)$ and pass each individual $\mathbb{O}_{j}\in\mathbb{R}^{M\times D}$ through a $\mathrm{LayerNorm}$ layer.
IV-F Multi-scale Wavelet Predictor
@@
In order to reconstruct the predicted series with length $H$, the length of high-pass and low-pass coefficients of each decomposition layers input to the IDWT should adhere to the same rules as the DWT process. Specifically, for $\hat{y}_{h}^{(j)}\in\mathbb{R}^{M\times H^{(j)}}$ and $\hat{y}_{l}^{(J)}\in\mathbb{R}^{M\times H^{(J)}}$, we set $H^{(j)}=\lfloor\frac{H^{(j-1)}+S-1}{2}\rfloor$ and $H^{(0)}=H$. To map the embeddings to the specific $H^{(j)}$.
For the output $\mathbb{O}_{m}^{(N)}$ of the last RoRA layer, we transpose it to $\mathbb{O}_{j}\in\mathbb{R}^{M\times D}$ where $j\in\{1,2,...,J+1\}$ and pass them through a corresponding MLP individually, as defined in Eq 10:
$\widehat{\mathcal{C}}$ $=\mathrm{WavePredictor}_{j}(\mathbb{O}_{j})$ (10)
@@
where $\widehat{\mathcal{C}}=\{\hat{yh}^{(1)},\hat{yh}^{(2)}...,\hat{yh}^{(J)},\hat{yl}^{(J)}\}$, $W_{j}\in\mathbb{R}^{H^{(j)}\times D}$ and $b_{j}\in\mathbb{R}^{H^{(j)}\times 1}$ for $j\in\{1,2,...,J\}$, while $W_{J+1}\in\mathbb{R}^{H^{(J)}\times D}$ and $b_{J+1}\in\mathbb{R}^{H^{(J)}\times 1}$. We then input the obtained predicted high-pass and low-pass coefficients to IDWT to get the final outputs $X_{out}\in\mathbb{R}^{H\times M}$.
@@ V-A Datasets
We evaluate our model on 8 real-world datasets: Weather, Electricity, Traffic, Solar and 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2). These datasets are extensively utilized, covering multiple fields including weather, energy management and transportation. The statistics of the datasets are shown in Table I.
@@
Dataset Variables Frequency Length
Weather 21 10 min 52696
Electricity 321 1 hour 26304
Traffic 862 1 hour 17544
ETTh1 7 1 hour 17420
ETTh2 7 1 hour 17420
ETTm1 7 15 min 69680
ETTm2 7 15 min 69680
Solar 137 10min 52179
V-B Baselines
V-C Experiment Settings
We set a uniform input length $L=96$ for all models and conduct experiments with $H=\{96,192,336,720\}$. By default, WaveRoRA employs a 4-level DWT using the Symlet3[25] wavelet basis. We map the wavelet coefficients to embeddings with $D=128$. For datasets with more variables, such as Weather, Traffic, Electricity and Solar, we construct 3-layer RoRA network, while for datasets with less variables, such as ETT datasets, we set RoRA layer number $N$ to 2. For RoRA, we set the token number of $R$ to $r=\min(20,\log{M})$. The attention heads number is set to 8 by default.
We replicate and run PatchTST, iTransformer and DLinear using the publicly available optimal parameters. For other baseline models, we adhere to their established implementations. We record MSE and MAE to measure the performance of different models. All experiments are conducted on one NVIDIA A800 GPU with 40GB memory to exclude the influence of platforms.
@@ V-C Experiment Settings
192 0.357 0.387 0.385 0.401 0.383 0.396 0.877 0.656 0.426 0.434 0.473 0.472 0.402 0.414
336 0.396 0.418 0.425 0.433 0.423 0.431 1.043 0.731 0.477 0.479 0.584 0.535 0.452 0.452
720 0.411 0.437 0.434 0.449 0.424 0.442 1.104 0.763 0.453 0.490 0.795 0.642 0.462 0.468
ETTm1 96 0.324 0.362 0.339 0.374 0.326 0.363 0.404 0.426 0.510 0.492 0.344 0.372 0.338 0.375
192 0.367 0.386 0.380 0.394 0.362 0.385 0.450 0.451 0.514 0.495 0.383 0.393 0.374 0.387
336 0.400 0.409 0.417 0.418 0.393 0.409 0.532 0.515 0.521 0.497 0.414 0.416 0.410 0.411
720 0.466 0.444 0.478 0.452 0.451 0.439 0.666 0.589 0.537 0.513 0.474 0.452 0.478 0.450
ETTm2 96 0.172 0.255 0.189 0.274 0.177 0.261 0.287 0.366 0.205 0.293 0.190 0.287 0.187 0.267
192 0.240 0.301 0.254 0.313 0.247 0.309 0.414 0.492 0.278 0.336 0.274 0.349 0.249 0.309
336 0.301 0.339 0.315 0.352 0.312 0.351 0.597 0.542 0.343 0.379 0.376 0.421 0.321 0.351
720 0.398 0.396 0.413 0.405 0.424 0.416 1.073 1.042 0.414 0.419 0.545 0.519 0.408 0.403
Solar 96 0.197 0.215 0.207 0.241 0.202 0.240 0.310 0.331 0.864 0.707 0.284 0.372 0.250 0.292
192 0.231 0.252 0.240 0.264 0.240 0.265 0.734 0.725 0.844 0.695 0.320 0.396 0.296 0.318
336 0.248 0.273 0.251 0.274 0.249 0.276 0.750 0.735 0.751 0.700 0.354 0.421 0.319 0.330
720 0.248 0.274 0.253 0.277 0.248 0.277 0.769 0.765 0.867 0.710 0.354 0.415 0.338 0.337
TABLE III: Ablation results of MTS forecasting task on 8 real-world datasets. Each mse and mae score is the average value of $H=\{96,192,336,720\}$. The best results are in bold.

### v2

@@ IV-C Multi-scale Wave Embedding
The input MTS data is first decomposed to $J$ high-pass coefficients $y_{h}^{(j)}\in\mathbb{R}^{M\times L^{(j)}}$ and 1 low-pass coefficients $y_{l}^{(J)}\in\mathbb{R}^{M\times L^{(J)}}$ with $J$-level DWT, where $j\in\{1,2,...,J\}$. Each coefficient series is passed through a corresponding wave-embedding layer defined in Eq 7, where $\mathcal{C}=\{y_{h}^{(1)},y_{h}^{(2)}...,y_{h}^{(J)},y_{l}^{(J)}\}$ and $\mathrm{WaveEmbedding}$ is a linear projector. Specifically, $\mathrm{WaveEmbedding}_{j}$ contains a weight matrix $W_{j}$ and a bias matrix $b_{j}$. Note that $W_{j}\in\mathbb{R}^{D\times L^{(j)}}$ for $j\in\{1,2,...,J\}$ and $W_{J+1}\in\mathbb{R}^{D\times L^{(J)}}$, while $b_{j}\in\mathbb{R}^{D\times 1}$. The multi-scale wave-embedding layers unify the feature dimensions of wavelet coefficients obtained from the multi-layer DWT. What’s more, the embedding process facilitates in capturing the temporal patterns within the time series.
@@
$\mathcal{M}_{\Theta,n}^{d}=\begin{pmatrix}\cos{n\theta_{1}}&-\sin{n\theta_{1}}&0&0&\cdots&0&0\\ \sin{n\theta_{1}}&\cos{n\theta_{1}}&0&0&\cdots&0&0\\ 0&0&\cos{n\theta_{2}}&-\sin{n\theta_{2}}&\cdots&0&0\\ 0&0&\sin{n\theta_{2}}&\cos{n\theta_{2}}&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&0&\cdots&\cos{n\theta_{\frac{d}{2}}}&-\sin{n\theta_{\frac{d}{2}}}\\ 0&0&0&0&\cdots&\sin{n\theta_{\frac{d}{2}}}&\cos{n\theta_{\frac{d}{2}}}\end{pmatrix}$ (8)
IV-D WaveRoRA Encoder
RoRA is proposed to capture inter-series dependencies in the wavelet domain. We first transpose $\mathbb{E}\in\mathbb{R}^{(J+1)\times M\times D}$ into series-wise tokens $\mathbb{S}\in\mathbb{R}^{M\times((J+1)\cdot D)}$ where each token in $\mathbb{S}$ represents the concatenation of wavelet embeddings across all decomposition levels for a single variable. To simplify the representation, we use $D^{\prime}$ instead of $(J+1)\cdot D$ and $I\in\mathbb{R}^{M\times D^{\prime}}$ instead of $\mathbb{S}^{(n)}$ of each WaveRoRA Encoder.
IV-D1 Rotary Route Attention
A set of linear projectors are applied to generate query, key and value matrices, represented as $Q\in\mathbb{R}^{M\times D^{\prime}}$, $K\in\mathbb{R}^{M\times D^{\prime}}$ and $V\in\mathbb{R}^{M\times D^{\prime}}$. Subsequently, a small number of routing tokens $R\in\mathbb{R}^{r\times D^{\prime}}$ is randomly initialized and the multi-head version of RoRA can be represented in Alg 1, where $\odot$ refers to Hadamard product and the activation function $\sigma$ is set as $\mathrm{SiLU}$[37], generating $G\in\mathbb{R}^{M\times D^{\prime}}$ which serves as a gating function. The routing tokens $R$ is initially viewed as the query for $K$ and $V$, aggregating routing features from all keys and values. Afterward, $R$ and $V_{R}$ will serve as the key and value for $Q$ respectively, enabling the transfer of information from the routing features to each query token in $Q$. In our implementations, $R$ is expected to prune redundant relationships between low-correlation variables while preserving interaction information among high-correlation variables. In addition, we set $r\ll M$ as a fixed number to achieve linear complexity of $O(Mrd)$ relative to the variable number $M$.
RoRA could be simplified to a form of generalized linear attention as shown in Eq 10, which may lead to insufficient feature diversity[39]. Therefore, we try to solve this issue by injecting relative positional information through rotary mechanism. Typically, a rotary matrix $\mathcal{M}_{\Theta}^{d}$ defined in Eq 8 is multiplied with $Q$ and $K$ in SA or LA, where $\Theta$ represents a set of predefined rotational angles and $d$ is the dimension of each token. $\mathcal{M}_{\Theta}^{d}$ adaptively adjusts the angles based on each token’s position, thereby incorporating the relative positional information $(m-n)\Theta$ for any indexes $m$, $n$ in $Q$ and $K$. However, such mechanism relies on the assumption that $Q$ and $K$ are directly multiplied to generate the attention scores. In our proposed RoRA, the routing tokens $R$ may disrupt the rotational angles applied to $Q$ and $K$. Assuming that we follow the typical use of RoPE as shown in Eq 9, the relative positional information would not be correctly computed. Specifically, the hypothetical rotary attention can be expressed as Eq 10, where $\Psi$ could be a certain kernel function. Assuming that $\Psi$ satisfies linear properties, we can then get an approximate solution $\Psi(Q\mathcal{M}^{D^{\prime}}_{\Theta}R^{T}R(\mathcal{M}^{D^{\prime}}_{\Theta})^{T}K^{T})$. Furthermore, it can easily be demonstrated that for any position indexes $m$ and $n$, the relative positional embeddings with the rotational angle $(m-n)\Theta$ could only be correctly obtained when $R^{T}R$ equals to an identity matrix. However, the routing tokens are randomly initialized and $\Psi$ could also be non-linear functions like Softmax. Therefore, applying rotary matrix to $Q$ and $K$ may not properly model the relative positional information.
$V_{R}^{\prime}$ $=\mathrm{Attention}(R,K\mathcal{M}^{D^{\prime}}_{\Theta},V)$ (9)
$V_{O}^{\prime}$ $=\mathrm{Attention}(Q\mathcal{M}^{D^{\prime}}_{\Theta},R,V_{R})$
$V_{O}^{\prime}$ $=\Psi(Q\mathcal{M}^{D^{\prime}}_{\Theta}R^{T})\Psi(R(K\mathcal{M}^{D^{\prime}}_{\Theta})^{T})V$ (10)
$\approx\Psi\big{(}Q\mathcal{M}^{D^{\prime}}_{\Theta}R^{T}R(\mathcal{M}^{D^{\prime}}_{\Theta})^{T}K^{T}\big{)}V$
@@ IV-D1 Rotary Route Attention
Algorithm 1 The implementation of multi-head RoRA
0:  Input sequence tokens $I\in\mathbb{R}^{M\times D^{\prime}}$
0:  Output representations $O\in\mathbb{R}^{M\times D^{\prime}}$
1:  $G\in\mathbb{R}^{N\times D^{\prime}}\leftarrow\sigma\big{(}\mathrm{Linear}(I)\big{)}$
2:  $R\in\mathbb{R}^{r\times D^{\prime}}\leftarrow$ Randomly initialized
3:  for $h=1$ to $\mathrm{H}$ do
4:     $Q^{h}\in\mathbb{R}^{M\times(D^{\prime}/H)}\leftarrow\mathrm{Linear}(I)$
5:     $K^{h}\in\mathbb{R}^{M\times(D^{\prime}/H)}\leftarrow\mathrm{Linear}(I)$
6:     $V^{h}\in\mathbb{R}^{M\times(D^{\prime}/H)}\leftarrow\mathrm{Linear}(I)$
7:     $R^{h}\in\mathbb{R}^{r\times(D^{\prime}/H)}\leftarrow\mathrm{Linear}(R)$
8:     $A_{R-K}^{h}\in\mathbb{R}^{r\times M}\leftarrow\big{(}\Psi(RK^{T})\mathcal{M}^{r}_{\Theta}\big{)}^{T}$
9:     $V_{R}^{h}\in\mathbb{R}^{r\times(D^{\prime}/H)}\leftarrow A_{R-K}^{h}V^{h}$
10:     $A_{Q-R}^{h}\in\mathbb{R}^{M\times r}\leftarrow\Psi(QR^{T})\mathcal{M}^{r}_{\Theta}$
11:     $V_{O}^{h}\in\mathbb{R}^{M\times(D^{\prime}/H)}\leftarrow A_{Q-R}^{h}V_{R}^{h}+\mathrm{Linear}(V^{h})$
12:  end for
13:  $O\in\mathbb{R}^{M\times D^{\prime}}\leftarrow\mathrm{Linear}\big{(}\mathrm{Concat}(V_{O}^{1},V_{O}^{2},...,V_{O}^{H})\odot G\big{)}$
14:  return $O$
Based on the conclusions, we apply $\mathcal{M}^{r}_{\Theta}$ to the attention scores of $Q$-$R$ and $R$-$K$ pairs, as shown in Alg 1. The basic formula of RoRA can then be written as Eq 11, where $\Psi$ here is the Softmax function. In this case, the rotary matrix properly reflect the relative positional differences between elements. Although the corresponding dimensionality is reduced to $r$, RoRA could take advantages of the nonlinearity of Softmax and the mapping mechanism of $R$. In addition, we introduce a lienar projector as the skip-connection module to further enhance the feature diversity of $Q$ and $K$.
$V_{O}=$ $\Psi(QR^{T})\mathcal{M}^{r}_{\Theta}(\mathcal{M}^{r}_{\Theta})^{T}\Psi(RK^{T})^{T}V+\mathrm{Linear}(V)$ (11)
IV-D2 Wavelet-wise Normalization
@@
Considering that the wavelet coefficients of different decomposition layers correspond to different frequency intervals, we transpose the output of RoRA to wavelet-wise tokens again and apply wavelet-wise normalization. The process is defined in Eq 12, where $\mathrm{RoRA}$ corresponds to Alg 1 , $\mathbb{S}^{(n)}\in\mathbb{R}^{M\times((J+1)\cdot D)}$ is the series-wise tokens, $\mathbb{W}^{(n)}\in\mathbb{R}^{(J+1)\times M\times D}$ is the wavelet-wise tokens and $n$ represents the number of WaveRoRA Encoder layer.
$\mathbb{W}^{(n)}$ $=\mathrm{Transpose}(\mathrm{RoRA}(\mathbb{S}^{(n)}))$ (12)
$\mathbb{S}^{(n+1)}$ $=\mathrm{Transpose}(\mathrm{WaveNorm}(\mathbb{W}^{(n)}))$
IV-E Multi-scale Wavelet Predictor
@@
In order to reconstruct the predicted series with length $H$, the length of high-pass and low-pass coefficients of each decomposition layers input to the IDWT should adhere to the same rules as DWT. Specifically, for $\hat{y}_{h}^{(j)}\in\mathbb{R}^{M\times H^{(j)}}$ and $\hat{y}_{l}^{(J)}\in\mathbb{R}^{M\times H^{(J)}}$, we set $H^{(j)}=\lfloor\frac{H^{(j-1)}+S-1}{2}\rfloor$ and $H^{(0)}=H$. To map the embeddings to the specific $H^{(j)}$, we transpose the output of the last RoRA layer to wavelet embeddings $\mathbb{O}\in\mathbb{R}^{(J+1)\times M\times D}$ and pass them through a corresponding MLP individually, defined as Eq 13:
$\widehat{\mathcal{C}}_{j}$ $=\mathrm{WavePredictor}_{j}(\mathbb{O}_{j})$ (13)
@@
where $\widehat{\mathcal{C}}=\{\hat{y_{h}}^{(1)},\hat{y_{h}}^{(2)}...,\hat{y_{h}}^{(J)},\hat{y_{l}}^{(J)}\}$ contains the predicted high-pass and low-pass coefficients, $W_{j}\in\mathbb{R}^{H^{(j)}\times D}$ and $b_{j}\in\mathbb{R}^{H^{(j)}\times 1}$ for $j\in\{1,2,...,J\}$, while $W_{J+1}\in\mathbb{R}^{H^{(J)}\times D}$ and $b_{J+1}\in\mathbb{R}^{H^{(J)}\times 1}$. We set $\widehat{\mathcal{C}}$ as the input of IDWT to get the final outputs $X_{out}\in\mathbb{R}^{H\times M}$.
@@ V-A Datasets
We evaluate our model on 8 real-world datasets: Weather, Electricity, Traffic, Solar and 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2)[12]. These datasets are extensively utilized, covering multiple fields including weather, energy management and transportation. The statistics of the datasets are shown in Table I.
@@
Dataset Weather Electricity Traffic ETTh1 ETTh2 ETTm1 ETTm2 Solar
Variables 21 321 862 7 7 7 7 137
Frequency 10 min 1 hour 1 hour 1 hour 1 hour 15 min 15 min 10 min
@@
192 0.356 0.387 0.385 0.401 0.383 0.396 0.877 0.656 0.426 0.434 0.473 0.472 0.402 0.414
336 0.394 0.416 0.425 0.433 0.423 0.431 1.043 0.731 0.477 0.479 0.584 0.535 0.452 0.452
720 0.403 0.433 0.434 0.449 0.424 0.442 1.104 0.763 0.453 0.490 0.795 0.642 0.462 0.468
ETTm1 96 0.320 0.357 0.339 0.374 0.326 0.363 0.404 0.426 0.510 0.492 0.344 0.372 0.338 0.375
192 0.365 0.383 0.380 0.394 0.362 0.385 0.450 0.451 0.514 0.495 0.383 0.393 0.374 0.387
336 0.396 0.404 0.417 0.418 0.393 0.409 0.532 0.515 0.521 0.497 0.414 0.416 0.410 0.411
720 0.463 0.442 0.478 0.452 0.451 0.439 0.666 0.589 0.537 0.513 0.474 0.452 0.478 0.450
ETTm2 96 0.175 0.259 0.189 0.274 0.177 0.261 0.287 0.366 0.205 0.293 0.190 0.287 0.187 0.267
192 0.241 0.302 0.254 0.313 0.247 0.309 0.414 0.492 0.278 0.336 0.274 0.349 0.249 0.309
336 0.304 0.342 0.315 0.352 0.312 0.351 0.597 0.542 0.343 0.379 0.376 0.421 0.321 0.351
720 0.405 0.398 0.413 0.405 0.424 0.416 1.073 1.042 0.414 0.419 0.545 0.519 0.408 0.403
Solar 96 0.197 0.222 0.207 0.241 0.202 0.240 0.310 0.331 0.864 0.707 0.284 0.372 0.250 0.292
192 0.232 0.254 0.240 0.264 0.240 0.265 0.734 0.725 0.844 0.695 0.320 0.396 0.296 0.318
336 0.250 0.275 0.251 0.274 0.249 0.276 0.750 0.735 0.751 0.700 0.354 0.421 0.319 0.330
720 0.247 0.274 0.253 0.277 0.248 0.277 0.769 0.765 0.867 0.710 0.354 0.415 0.338 0.337
V-B Baselines
V-C Experiment Settings
We set a uniform input length $L=96$ for all models and conduct experiments with $H=\{96,192,336,720\}$. By default, WaveRoRA employs a 4-level DWT using the Symlet3[22] wavelet basis. We map the wavelet coefficients to embeddings with $D=64$. For datasets with less variables, such as ETT datasets, we set RoRA layer number $N$ to 2; while for those with more variables like Weather, Electricity and Solar, we set $N=3$; For Traffic that has large number of variables, we set $N=4$. For RoRA, we set the token number of $R$ to $r=\min(10,\frac{\log{M}+\sqrt{M}}{2})$. The attention heads number is set to 8 by default.
We replicate and run PatchTST, iTransformer and DLinear using the publicly available optimal parameters. For other baseline models, we adhere to their established implementations. We record the MSE and MAE to measure the performance of different models. These metrics are defined in Eq 14 and Eq 15, where $Y$ and $\hat{Y}$ represent ground truth and predicted values respectively. All experiments are conducted on one NVIDIA A800 GPU with 40GB memory to exclude the influence of platforms.
$\mathrm{MSE}(Y,\hat{Y})$ $=\frac{1}{|Y|}\sum_{i=1}^{|Y|}|Y_{i}-\hat{Y}_{i}|^{2}$ (14)
$\mathrm{MAE}(Y,\hat{Y})$ $=\frac{1}{|Y|}\sum_{i=1}^{|Y|}|Y_{i}-\hat{Y}_{i}|$ (15)
TABLE III: Ablation results of MTS forecasting task on Weather, Traffic and Electricity. Each MSE and MAE score is the average value of $H=\{96,192,336,720\}$. The best results are in bold.